RESONANCE ENHANCED TWO-PHOTON IONIZATION (RE2PI) SPECTRUM OF THE 520 NM SYSTEM OF RbCs

\underline{YOUNGJEE YOON}, YONGHOON LEE, SUN JONG BAEK, JEONGYE CHOI, and BONGSOO KIM, Department of Chemistry, Korea Advanced Institute of Science and Technology, Taejon 305-701, Korea.

The rovibrational spectrum of RbCs molecule in the range of 520 nm is observed by the resonance enhanced two-photon ionization (RE2PI) method. A very cold pulsed molecular beam that contains RbCs, Rb_{2}, Cs_{2}, etc is generated by a high temperature pulsed nozzle. Only RbCs^{+} ion could be detected using the time-of-flight (TOF) mass spectrometer (mass resolution ∼ 1300). Vibrational bands (v′ = 6 ∼ 31) are rotationally resolved by the high resolution dye laser with an intracavity etalon (laser resolution ∼ 0.02 cm^{−1}). The excited electronic state is assigned to the 5 ^{1}Σ^{+} state that dissociates into Rb(5s ^{2}S_{1/2}) + Cs(7s ^{2}S_{1/2}). By the pseudopotential calculation, the 5 ^{1}Σ^{+} state has adiabatic potential curve with a shelf as a result of the avoided crossing with ionic pair state at long internuclear distances. Observed in the Franck-Condon region, however, the ΔG_{v} curve shows slightly positive curvature for v′ > 20, which may result from the avoided crossing with the 6 ^{1}Σ^{+} state at ∼ 6 Å. By the selection rules of ΔJ = ±1 for ^{1}Σ → ^{1}Σ transitions, only P and R lines are observed. The rotational constants, B_{v}, and the vibrational term value, T_{v}, are determined from the analysis of the rotationally resolved spectra. From the vibrational energy level spacing, ΔG_{v}, the spectroscopic parameters for the 5^{1}Σ^{+} state are determined as T_{e} = 18560.12(8) cm^{−1}, ω_{e} = 40.83(1) cm^{−1}, and ω_{e}x_{e} = 0.2465(6) cm^{−1}. The dissociation energy, D_{e}, is 3811.5 cm^{−1}. Measured isotope shifts confirm the absolute vibrational numbering. The B_{e} and α_{e} obtained from the B_{v} vs. v plot are 0.013468(8) cm^{−1} and 7.48(4) × 10^{−5} cm^{−1}, respectively. Using these molecular constants, potential energy curve of the 5 ^{1}Σ^{+} state is constructed by the RKR method.